OpenStudy (anonymous):

Without graphing, determine whether the graphs of the equations are identical lines, parallel lines, or lines intersecting at a single point. Choose one answer. a. parallel lines b. identical lines c. lines intersecting at a single point

OpenStudy (anonymous):

can you post a screenshot?

OpenStudy (ash2326):

yeah attachement is directing to a login page

OpenStudy (anonymous):

Sorry! =)

OpenStudy (anonymous):

You can re-arrange and compare slopes e.g. 6x - y = 25 --> y = 6x-25 x + 5y =30 --> y =-(1/5)x+6 The slopes are different so the lines will intersect.

OpenStudy (anonymous):

\[(1) = 6x-y=25\]\[(2)= x+5y=30\] Multiplying (1) by 5 and adding it with (2), we get \[5 \times (6x-y)+x+5y=5 \times 25 + 30\]\[30x - 5y + x + 5y = 125 + 30\]\[31x=155\]\[x=155/31\]\[y=6 \times 155/31 - 25\] So we now know that the lines interset and one and only point.